# Hierarchical Balance Packing: Towards Efficient Supervised Fine-tuning for Long-Context LLM

### v1

@@ Abstract
Training Long-Context Large Language Models (LLMs) is challenging, as hybrid training with long-context and short-context data often leads to workload imbalances. Existing works mainly use data packing to alleviate this issue but fail to consider imbalanced attention computation and wasted communication overhead. This paper proposes Hierarchical Balance Packing (HBP), which designs a novel batch-construction method and training recipe to address those inefficiencies. In particular, the HBP constructs multi-level data packing groups, each optimized with a distinct packing length. It assigns training samples to their optimal groups and configures each group with the most effective settings, including sequential parallelism degree and gradient checkpointing configuration. To effectively utilize multi-level groups of data, we design a dynamic training pipeline specifically tailored to HBP, including curriculum learning, adaptive sequential parallelism, and stable loss. Our extensive experiments demonstrate that our method significantly reduces training time over multiple datasets and open-source models while maintaining strong performance. For the largest DeepSeek-V2 (236B) MOE model, our method speeds up the training by 2.4 with competitive performance.
@@ 1 Introduction
Large Language Models (LLMs) (Dubey et al., 2024; Yang et al., 2024; Cai et al., 2024) have achieved state-of-the-art performance in tasks like machine translation, summarization, and code generation. However, many applications demand to process and understand long-context information (Chen et al., 2023a; Peng et al., 2023; Bai et al., 2024a), such as summarizing books, analyzing legal documents, or retaining context in multi-turn conversations. This underscores the necessity for long-context LLMs that can efficiently process long input sequences.
@@
As mentioned in (Dubey et al., 2024), incorporating long context during the Supervised Fine-Tuning (SFT) (Ouyang et al., 2022) stage is highly necessary. On the other hand, general short-context data is also crucial to maintain the model’s general capabilities. However, this hybrid dataset composition introduces significant challenges, primarily in terms of speed and accuracy. For speed, long-context data intensifies training inefficiencies due to imbalanced workloads; For accuracy, long-context data can degrade performance on short-context tasks, affecting the model’s general capabilities. These challenges hinder the efficiency and effectiveness of SFT for long-context LLMs.
@@
The workload imbalance caused by the hybrid of long and short data arises from two main aspects: (1) within mini-batch imbalance (Yao et al., 2024), which is caused by excessive padding from randomized mini-batch construction, and (2) across mini-batch imbalance (Yao et al., 2024), which is caused by uneven computation distribution over data parallel replicas. Existing approaches mainly address this issue using data packing (Bai et al., 2024a; Wang et al., 2024a), which combines variable-length data into fixed-length mini-batches. While data packing helps mitigate workload imbalances, it can also introduce new challenges. First, data packing alters the data distribution, which might affect the models’ performance. Second, the complexity of attention computation for short- and long-context data differs significantly. As illustrated in Figure 1, simply combining them has a high variance of data length, resulting in imbalanced attention computation and workload imbalance. Third, handling long-context data requires sequential parallelism (SP) and collective communication for attention computation (Jacobs et al., 2023; Liu & Abbeel, 2023), which short data do not need. When simply mixed together, short-context data must wait for long-context data, leading to wasted communication time. Larger models, such as DeepSeek-V2 (236B) (DeepSeek-AI, 2024) MOE (Mixture of Experts) models, introduce higher communication overhead due to the increased number of parameters.
@@
To overcome the limitations of data packing, we propose Hierarchical Balance Packing (HBP), an innovative method that proposes multi-level data packing instead of conventional single-level data packing. HBP consists of three key components: (1) what are the optimal packing groups? (2) How to assign training samples to their optimal group? (3) How to train long-context LLM with those data? Firstly, we propose hierarchical group auto-selection to determine the optimal packing-length group set and corresponding configurations, including the packing length, Gradient Checkpointing configuration (Li et al., 2014), and the SP degree (how many partitions the data is divided into). Secondly, we propose balance packing to allocate each sample to the optimal group, aiming to minimize imbalanced attention computation and communication overhead. Thirdly, we adopt alternative training between different packing groups, along with curriculum learning and a stable loss normalizer to stabilize the training process.
@@
We validate the effectiveness of our method through extensive experiments in multiple settings. For example, on datasets Tulu3 (Lambert et al., 2024a) (32K) + Longcite (Zhang et al., 2024) (128K), our approach speeds up 2.4 (57.1 to 23.8 GPU Days) on DeepSeek-V2 (236B) (DeepSeek-AI, 2024); On datasets OpenHerme (Teknium, 2023) (4K) + Longcite (128K), our approach reduces training time from 2.95 to 2.04 GPU Days about 1.45 speeds up on LLama3.1-8B (Dubey et al., 2024). More importantly, our method preserves performance on both short- and long-context datasets while achieving significant efficiency gains. Experiments on various models at different scales like LLama3.1-8B (Dubey et al., 2024), Qwen2.5-32B (Yang et al., 2024), Qwen2.5-72B (Yang et al., 2024), and DeepSeek-V2 (236B) demonstrate consistent improvements, showing the effectiveness and generalizability of our approach.
@@ 2.1 Long-Context LLM
Long-Context Extension. Long-Context Extensions aim to enhance LLMs’ capabilities in handling long contexts. Current research can be broadly categorized into approaches that require fine-tuning and those that operate in a zero-shot manner. Zero-shot approaches often leverage techniques such as prompt compression (Jiang et al., 2023) or specially designed attention mechanisms (Sun et al., 2023; Han et al., 2024). On the other hand, fine-tuning methods primarily focus on extending position encoding, such as RoPE-based approaches (Peng et al., 2023), or utilizing memory-augmented architectures (Packer et al., 2023).
Long-Context Supervised Fine-Tuning. SFT aligns LLMs more effectively with user intent. For Long-Context SFT, research mainly concentrates on generating long-context datasets (Xiong et al., 2023) and establishing corresponding benchmarks (Chen et al., 2023b; Zhang et al., 2024; Bai et al., 2024b). In contrast, our work emphasizes training efficiency and performance. The most relevant work to ours is LongAlign (Bai et al., 2024a), which also focuses on issues related to workload balance and accuracy degradation. They proposed using packing and loss-reweighting to mitigate these issues. However, they failed to recognize the imbalanced attention computation and wasted communicated overhead due to the packing of short- and long-context data, which limits efficiency.
@@ 2.2 Data Packing
Data packing (Wang et al., 2024a) is a more practical approach compared to randomly organizing data batches in LLM training. It reduces padding within batches and minimizes idle time across different data-parallel groups. Common packing methods include Random Packing (Contributors, 2023b), Sorted Batching (Kundu et al., 2024), First Fit Shuffle (FFS), First Fit Decrease (FFD), Best Fit Shuffle (BFS), Shortest-Pack-First Histogram-Packing (SPFHP) (Krell et al., 2021), Iterative sampling and filtering (ISF) (Yao et al., 2024). However, all those packing methods operate on a fixed length. In contrast, our method introduces multiple packing groups with varying lengths, which enables more flexible handling of hybrid training involving short- and long-context data.
@@ 3.1 Measuring Metrics
Dist Balance Ratio (DBR) (Yao et al., 2024) quantifies the computational balance inter-devices based on input length.
Padding Ratio (PR) (Yao et al., 2024) measures the proportion of wasted computations resulting in intra-device from padding based on input length.
Attention Balance Ratio (ABR) is proposed to quantify the imbalance with respect to attention computation for different data inter-devices.
Previous metrics estimate the computational cost of attention based solely on the length of the input with full attention. However, using packing algorithms, attention computation becomes a significant factor in the overall cost. Consider the packing of a sequence as an example, both and have the same total length. However, their actual attention computation differs significantly, with complexities of and , respectively. The Attention Balance Ratio (ABR) is given by:
The cost of attention is proportional to the attention balance ratio, Cost(Attn) ABR.
Communication Ratio (CR) is proposed to measure the additional communication overhead.
Sequence parallelism (SP) is crucial in long-context LLM training. However, this introduces the trade-off of increased communication overhead. The additional communication during hybrid training depends on the input sample length. For instance, an 8k input may not require any SP partitions, while a 32k input might necessitate multiple SP partitions, significantly increasing communication costs.
Average Tokens (Ave-T) is defined as the average number of tokens processed per iteration, serving as a measure of the model’s workload.
A small token count per iteration indicates computational inefficiency. Increasing the batch size appropriately can improve Ave-T throughput and overall training efficiency.
@@ 3.2 Packing Analysis
The packing strategy is widely utilized in SFT to improve training efficiency. We first conducted a comprehensive analysis to validate its effectiveness. However, our study also revealed several challenges and limitations associated with this line of approaches.
Importance of Packing. In Table 2, we conduct three batching alternatives, random batching, ISF packing batching (Yao et al., 2024) (comparison between different packing methods is shown in Table 8), sorted batching (Bai et al., 2024a) (ensures that the sequences within each batch have similar lengths) at three sequence length 4K, 32K, 128K using Tulu3 (Lambert et al., 2024a) dataset. Both the sorted and packing strategies significantly reduce DBR and PR, leading to substantial improvements in training speed. However, in longer sequence scenarios such as 128K, the packing strategy has a higher average of tokens (Ave-T) compared to the sorted method, achieving a higher GPU utilization. This underscores the importance of the packing strategy for training on mixed-length datasets.
@@
| 128K | sorted | 0 | 0.0 | 0.9K | 33.3 (1.1) |
| 128K | packing | 0.001 | 0.0 | 128K | 5.2 (7.3) |
Limitation of Packing. Although packing can partially address the efficiency issues associated with hybrid training, several limitations remain. Specifically, samples with varying sequence lengths exhibit different complexities in attention computation. Directly mixing these samples can result in a workload imbalance. As illustrated in Table 3, the ABR increases significantly with the growth of sequence length, which indicates a rise in device idle time. Moreover, long sequences necessitate communication for attention computation, while short sequences do not. Directly mixing them can lead to extra communication overhead. As shown in Table 3, the CR reaches 1 when the sequence length increases, indicating that all short-context data are involved in unnecessary communication processes.
| Packing Len | SP | ABR | CR | DBR | PR |
| 4K | 1 | 0.434 | 0 | 0.003 | 0.0 |
| 32K | 4 | 0.616 | 1.0 | 0.001 | 0.0 |
| 128K | 8 | 0.667 | 1.0 | 0.003 | 0.0 |
3.3 Training Strategy Analysis
@@ 4 Hierarchical Balance Packing
In this section, we first introduce how to determine the optimal packing length groups in Section 4.1. Then, we present how to fit each sample to the best packing group in Section 4.2. At last, we design a tailored dynamic training pipeline for HBP in Section 4.3. The overall framework is illustrated in Figure 2.
@@ 4.1 Hierarchical Groups Auto-Selection
To determine the optimal packing length groups, we design a profile-based auto-selection algorithm as described in Algorithm 1. It operates in two stages: (1) find the best training strategy for predefining possible sequence length set (e.g., 8K, 16K, 32K, 64K, 128K) based on naive packing. (2) deriving the final packing groups by optimizing communication overhead.
Stage 1: Find the best training strategy. The algorithm begins by initializing two empty lists, and , to store the profiling times and the corresponding strategies, respectively. For each possible input length , we compute the optimal SP degree and GC configuration with the FindBestSpCkpt function, which achieves the optimal trade-off between SP degree and GC configuration.
Specifically, given a packing length , we iterate all possible degrees and GC configurations and profile their iteration time. The best combination is then found by a binary search or greedy method. More details on the implementation can be found in Appendix B. Once all input lengths are processed, the algorithm selects the best strategy by identifying the index that minimizes the profiling times in . The optimal strategy and input length are:
@@
After calculating the optimal and the corresponding , we can optimize it further. The can an individual sample to be split across different SP processes, leading to additional communication overhead during attention computation. To mitigate this, we propose a method that ensures smaller samples remain intact during SP training. When and , instead of directly packing data into 32K sequences, we first partition into smaller chunks of . Data with lengths smaller than 8K are packed into these 8K chunks first, which are then combined to form 32K sequences. By structuring the data this way, SP training can proceed without incurring any communication overhead from attention computation.
For , the smallest packing unit is derived based on its optimal SP degree. Similarly, the smallest packing unit for is also considered. If is smaller than , it will be merged into the range. Finally, the hierarchical packing groups is obtained.
@@ 4.2 Balance Packing
After obtaining the optimal hierarchical pack groups, we distribute the dataset samples into different groups while ensuring that the metrics outlined in Section 3.1 (DBR, PR, ABR, and CR) are well-optimized, of which conventional packing struggles. We first divide the entire dataset into sub-datasets based on . For each , the following steps are executed:
@@
(2) GreedyFill: We use the remaining unpacked data to perform a greedy fill since larger groups find it difficult to fill the packed data with only their own data.
(3) Sorting and Batching: We sort elements in based on attention complexity and construct mini-batches according to global token number requirements.
@@
The procedure of balance packing is illustrated in Algorithm 2. Since our approach inherently involves multiple levels, i.e., hierarchical packing groups, it automatically separates short- and long-context data, avoiding wasted communication overhead and imbalanced attention computation and reducing CR and ABR significantly. We also achieve extremely low PR and DBR at multiple levels, thanks to GreedyFill. More Details about GroupData, GreedyFill, and Sorting are shown in Appendix C.

### v3

@@ Abstract
Training Long-Context Large Language Models (LLMs) is challenging, as hybrid training with long-context and short-context data often leads to workload imbalances. Existing works mainly use data packing to alleviate this issue, but fail to consider imbalanced attention computation and wasted communication overhead. This paper proposes Hierarchical Balance Packing (HBP), which designs a novel batch-construction method and training recipe to address those inefficiencies. In particular, the HBP constructs multi-level data packing groups, each optimized with a distinct packing length. It assigns training samples to their optimal groups and configures each group with the most effective settings, including sequential parallelism degree and gradient checkpointing configuration. To effectively utilize multi-level groups of data, we design a dynamic training pipeline specifically tailored to HBP, including curriculum learning, adaptive sequential parallelism, and stable loss. Our extensive experiments demonstrate that our method significantly reduces training time over multiple datasets and open-source models while maintaining strong performance. For the largest DeepSeek-V2 (236B) MoE model, our method speeds up the training by 2.4 with competitive performance. Codes will be released at https://github.com/ModelTC/HBP.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3] have achieved state-of-the-art performance in tasks like machine translation, summarization, and code generation. However, many applications demand to process and understand long-context information [4, 5, 6], such as summarizing books, analyzing legal documents, or retaining context in multi-turn conversations. This underscores the necessity for long-context LLMs that can efficiently process long input sequences.
@@
As mentioned in [1], incorporating long context during the Supervised Fine-Tuning (SFT) [7] stage is highly necessary. On the other hand, general short-context data is also crucial to maintain the model’s general capabilities. However, this hybrid dataset composition introduces significant challenges, primarily in terms of speed and accuracy. For speed, long-context data intensifies training inefficiencies due to imbalanced workloads; For accuracy, long-context data can degrade performance on short-context tasks, affecting the model’s general capabilities. These challenges hinder the efficiency and effectiveness of SFT for long-context LLMs.
@@
The workload imbalance caused by the hybrid of long and short data arises from two main aspects: (1) within mini-batch imbalance [8], which is caused by excessive padding from randomized mini-batch construction, and (2) across mini-batch imbalance [8], which is caused by uneven computation distribution over data parallel replicas. Existing approaches mainly address this issue using data packing [6, 9], which combines variable-length data into fixed-length mini-batches. While data packing helps mitigate workload imbalances, it can also introduce new challenges. First, data packing alters the data distribution, which might affect the models’ performance. Second, the complexity of attention computation for short- and long-context data differs significantly. The attention complexity of the packed samples is the sum of the attention complexities of all samples within each packing group. As illustrated in Figure 1, naive packing leads to high variance in attention complexity by directly mixing long and short samples, which causes imbalanced attention computation and workload imbalance. Third, handling long-context data requires sequential parallelism (SP) and collective communication for attention computation [10, 11], which short data do not need. When mixed, short-context data also requires sequence parallel communication, leading to wasted communication time. Larger models, such as DeepSeek-V2 (236B) [12] MoE (Mixture of Experts) models, introduce higher communication overhead due to the increased number of parameters.
@@
To overcome the limitations of data packing, we propose Hierarchical Balance Packing (HBP), an innovative method that proposes multi-level data packing instead of conventional single-level data packing. HBP consists of three key components: (1) What are the optimal packing groups? (2) How to assign training samples to their optimal group? (3) How can a long-context LLM be trained with that data? Firstly, we propose hierarchical group auto-selection to determine the optimal packing-length group set and corresponding configurations, including the packing length, Gradient Checkpointing configuration [13], and the SP degree (how many partitions the data is divided into). Secondly, we propose balanced packing to allocate each sample to the optimal group, aiming to minimize imbalanced attention computation and communication overhead. Thirdly, we adopt alternative training between different packing groups, along with curriculum learning and a stable loss normalizer to stabilize the training process.
@@
We validate the effectiveness of our method through extensive experiments in multiple settings. For example, on datasets Tulu3 [14] (32K) + Longcite [15] (128K), our approach speeds up 2.4 (57.1 to 23.8 GPU Days) on DeepSeek-V2 (236B) [12]; On datasets OpenHerme [16] (4K) + Longcite (128K), our approach reduces training time from 2.95 to 2.04 GPU Days about 1.45 speeds up on LLama3.1-8B [1]. More importantly, our method preserves performance on both short- and long-context datasets while achieving significant efficiency gains. Experiments on various models at different scales like LLama3.1-8B [1], Qwen2.5-32B [2], Qwen2.5-72B [2], and DeepSeek-V2 (236B) demonstrate consistent improvements, showing the effectiveness and generalizability of HBP.
@@ 2.1 Long-Context LLM
Long-Context Extension. Long-Context Extensions aim to enhance LLMs’ capabilities in handling long contexts. Current research can be broadly categorized into approaches that require fine-tuning and those that operate in a zero-shot manner. Zero-shot approaches often leverage techniques such as prompt compression [17] or specially designed attention mechanisms [18, 19]. On the other hand, fine-tuning methods primarily focus on extending position encoding, such as RoPE-based approaches [5], or utilizing memory-augmented architectures [20].
Long-Context Supervised Fine-Tuning. For Long-Context SFT, research mainly concentrates on generating long-context datasets [21] and establishing corresponding benchmarks [22, 15, 23]. LongAlign [6] also focuses on workload balance and accuracy degradation issues. LongAlign proposed using packing and loss-reweighting to mitigate these issues. However, they failed to recognize the imbalanced attention computation and wasted communication overhead due to the packing of short- and long-context data. FlexSP [24] dynamically organizes training samples into different data groups and uses flexible sequence parallelism to enable hybrid training. Its online data organization introduces significant overhead (5-15 seconds) each iteration and fails to handle attention computation complexity across data-parallel (DP) groups. In contrast, our method introduces only negligible overhead and achieves a better computation balance of attention.
@@ 2.2 Data Packing
Data packing [9] is a more practical approach compared to randomly organizing data batches in LLM training. It reduces padding within batches and minimizes idle time across different data-parallel groups. Common packing methods include Random Packing [25], Sorted Batching [6], First Fit Shuffle (FFS), First Fit Decrease (FFD), Best Fit Shuffle (BFS), Shortest-Pack-First Histogram-Packing (SPFHP) [26], Iterative sampling and filtering (ISF) [8]. However, those packing methods operate on a fixed length. HBP operates data globally by introducing multiple packing groups of varying lengths, enabling more flexible and efficient handling of hybrid training with both short- and long-context data.
@@ 3.1 Measuring Metrics
Dist Balance Ratio (DBR) [8] quantifies the computational balance inter-devices based on length. A lower DBR 1 indicates more balanced workload distribution across different devices.
Padding Ratio (PR) [8] measures the proportion of wasted computations resulting from intra-device padding based on input length. A lower PR 1 indicates fewer padding tokens.
Attention Balance Ratio (ABR) is proposed to quantify the imbalance in attention computation for different data inter-devices. Previous metrics estimate the computational cost of attention based solely on the length of the input with full attention. However, using packing algorithms, attention computation becomes a significant factor in the overall cost. Consider the packing of a sequence as an example, both and have the same total length. However, their actual attention computation differs significantly, with complexities of and , respectively. The Attention Balance Ratio (ABR) and example are given by 2. A lower ABR indicates a more balanced attention computation between different DP groups, resulting in less idle time.
Communication Ratio (CR) is proposed to measure the overhead of SP communication. in CR 3 represents the total number of tokens that require SP communication in the current iteration, while denotes the total number of tokens in the current iteration. A lower CR 3 indicates that fewer tokens require SP communication, resulting in reduced communication overhead.
Average Tokens (Ave-T) is defined as the average number of tokens processed per iteration, serving as a measure of the model’s workload. A larger Ave-T 3 indicates higher training efficiency.
@@ 3.2 Packing Analysis
Importance of Packing. In Table 4, we conduct three batching alternatives, random batching, ISF packing batching [8] (comparison between different packing methods is shown in Appendix A), sorted batching [6] (ensures that the sequences within each batch have similar lengths) at three sequence length 4K, 32K, 128K using Tulu3 [14] dataset. Both sorting and packing reduce DBR and PR, speeding up training. On longer sequences like 128K, packing achieves higher Ave-T than sorting, improving GPU utilization and benefiting mixed-length datasets.
Limitation of Packing. Although packing can partially address the efficiency issues associated with hybrid training, several limitations remain. Imbalance Problem of Attention Complexity: Packing samples with varying sequence lengths leads to differing attention computation complexities. Directly mixing these samples can cause imbalanced workloads and inefficient resource utilization. As illustrated in Table 2, the ABR increases significantly with the sequence length growth, indicating a rise in device idle time. More details are shown in Appendix D. SP Communication Overhead: long sequences require communication for attention computation, whereas short sequences do not. Directly mixing them can lead to extra communication overhead. As shown in Table 2, the CR reaches 1 when the sequence length increases, indicating that all short-context data are involved in unnecessary communication.
@@
| 128K | sorted | 0.01 | 0.02 | 125K | 5.5 (6.9) |
[…]
@@ 4 Hierarchical Balance Packing
In this section, we first describe how to determine optimal packing length groups (Section 4.1), then explain how each sample is assigned to a group (Section 4.2). Finally, we introduce a dynamic training pipeline for HBP (Section 4.3). The overall framework is shown in Figure 2.
@@ 4.1 Hierarchical Groups Auto-Selection
To determine the optimal packing length groups, we design a profile-based auto-selection algorithm as described in Algorithm 1. It operates in two stages: (1) finding the best training strategy for predefining a possible sequence length set (e.g., 8K, 16K, 32K, 64K, 128K) based on naive packing. (2) Deriving the final packing groups by optimizing communication overhead.
Stage 1: Find the best training strategy. The algorithm begins by initializing two empty lists, and , to store the profiling times and the corresponding strategies, respectively. For each possible input length , we compute the optimal degree of SP and the configuration of the GC with the FindBestSpCkpt function, which achieves the optimal trade-off between the degree of SP and the configuration of the GC. More details of FindBestSpCkpt are shown in Appendix I.1.
Specifically, given a packing length , we iterate all possible degrees and GC configurations and profile their iteration time. The best combination is then found by a greedy method. Once all input lengths are processed, the algorithm selects the best strategy by identifying the index that minimizes the profiling times in . The optimal strategy and input length are:
@@
After calculating the optimal and the corresponding , we can optimize it further to reduce communication overhead and obtain the final packing group lengths. Taking 32K packing groups in Figure 3 with SP=2 as an example, each SP process handles a split of 16K tokens. For samples longer than 16K, additional communication is required anyway. For samples shorter than 16K, we can group them into 16K packing groups, which is equivalent to training with 16K packing groups using SP=1, thereby reducing communication.
For , the smallest packing group is derived based on its optimal degree of SP , ensuring no communication overhead during sequential parallelism. Similarly, the smallest packing group for is also considered. If is smaller than , it will be merged into the range. Finally, the hierarchical packing groups is obtained. More implementation details are shown in Appendix I.
@@ 4.2 Balance Packing
After obtaining the optimal hierarchical pack groups, we distribute the dataset samples into different groups while ensuring that the metrics outlined in Section 3.1 (DBR, PR, ABR, and CR) are well-optimized, which conventional packing struggles with. We first divide the entire dataset into sub-datasets based on . For each , the following steps are executed:
@@
(2) GreedyFill: We use the remaining unpacked data for a greedy fill, as larger groups struggle to fill the packed data using only their samples. A simple example illustration is shown in Appendix E.
(3) Balance Batching: Constructing balance batches based on attention complexity using heuristic solution sorting support maintaining balanced attention computation across different packing groups. Heuristic solutions are an efficient, approximate solution for large-scale datasets. More evidences are shown in Appendix D.
@@
The procedure of balanced packing is illustrated in Algorithm 2. Since our approach inherently involves multiple levels, i.e., hierarchical packing groups, it automatically separates short- and long-context data, avoiding wasted communication overhead and imbalanced attention computation and reducing CR and ABR significantly. We also achieve low PR and DBR at multiple levels, thanks to GreedyFill. More details about the implementation are shown in Appendix J.
4.3 Dynamic Training Pipeline
Since HBP involves multi-level inputs, it is essential to design a dynamic training pipeline, enabling hot switching of different packing groups. Adaptive SP (pre-initialized multi-SP achieving zero overhead) is proposed to ensure efficient training with multi-level packing. Curriculum Learning and a Stable Loss Normalizer are proposed to improve the performance in hybrid training.
Adaptive Sequential Parallel: Each packing group is assigned an optimal training configuration , and we adopt an alternating scheme that selects the best SP degree and GC setting per packing group, as illustrated in Figure 2.
Curriculum Learning Strategy: Training on long-context tasks presents challenges because initiating training without instructional capabilities can result in significant fluctuations in the training loss, as illustrated in Figure 5. Thanks to our inherent hierarchical structure, it is straightforward to adopt a curriculum learning strategy that begins with general short-context data in the early stages of training. As training progresses, we transition to a hybrid approach that alternates training both short- and long-context data, as illustrated in Figure 2.
Stable Loss Normalizer: The training stability introduced by data packing is an important problem, as it impacts the data distribution. Previous work [6] has analyzed loss calculation, identifying two primary loss normalizers (Token-Mean) and (Sample-Mean):
where represents the local batch size within the DP group, and denotes the number of loss tokens in the batch . They argued that instability lies in the inconsistent loss of normalization. [27] also proposes sum loss without normalization to mitigate the effect of norm discrepancies. However, the sum loss introduces a trade-off: as the sequence length increases, the gradient values escalate disproportionately (1e+5), as shown in the left part of Figure 5. To address the above issues, we empirically observe (Average Token) of the global batch size across all DP groups can serve as a stable loss normalizer. The stable loss normalizer is derived from theoretical considerations, with the objective of guaranteeing that each token contributes equally to the aggregate loss. This formulation mitigates biases introduced by heterogeneous sequence lengths, varying data-parallel group sizes, and gradient accumulation strategies. We illustrate the effect of different normalization strategies under a simple setting: a data-parallel (DP) group of size , where each rank processes a local batch of samples.
Token-Mean: In the token-mean approach, the loss for each DP rank is normalized by its local token:
The global loss is then averaged across ranks:
This formulation normalizes losses rank-wise without accounting for global token distribution, potentially introducing bias when sequence lengths vary across ranks.
Sample-Mean: In the sample-mean method, losses are first normalized per sample and then averaged:
The global loss becomes:
This strategy balances individual samples but fails to reflect the total token count, giving undue weight to shorter sequences.
Stable Loss Normalizer: To eliminate such bias, we normalize by the global average token length:
Each DP’s loss is then normalized by , And the final loss is:
This normalization ensures that the contribution of each token is equal, regardless of which sample or rank it belongs to, thereby producing an unbiased global loss.
5 Experiments
5.1 Experimental Setup
Implementation Details. We use large-scale datasets: Tulu3 (32K)[14] for general tasks and LongCite (128K)[15] for long-context tasks. Both have shown strong performance across various benchmarks. They also enhance the model’s ability to handle input lengths from 0.1K to 128K tokens. We conduct experiments with the following models: LLaMA 3.1 [1], Qwen-2.5 [2], and DeepSeek-V2 (236B). Most models are trained on 32x H100 80GB GPUs using the DeepSpeed [28], while DeepSeek-V2 (236B) is trained with the Megatron-LM [29] with 256x H100 80G GPUs. In our experiments, we use DeepSpeed-Ulysses’s [10] sequence parallelism approach, and the ring-attention [11] method is also applicable. We conducted ablation experiments using the LLaMA3.1-8B model. For the Longsite dataset, approximately 2k samples are uniformly sampled. The loss normalizer for baselines without special instructions is Token-Mean, while HBP uses Ave-Token. GPU days are the evaluation metric to estimate the total training time. We use a learning rate of 1e-5, weight decay of 0.01, and adopt AdamW as our optimizer.
Evaluation. We comprehensively evaluated the LLM’s performance using OpenCompass [30]. For general tasks, several benchmark datasets were assessed, including MMLU [31], MMLU PRO [32], CMMLU [33], BBH [34], Math [35], GPQA Diamond [36], GSM8K [37], HellaSwag [38], MathBench [39], HumanEval [40], MBPP [41], IFEval [42], and Drop [43]. For long-context tasks, the evaluation including Ruler [44], NeedleBench [45], LongBench [46], and Longcite.
5.2 Main Results